### v1

@@ Abstract
Habitat assessment at local scales — critical for enhancing biodiversity and guiding conservation priorities — often relies on expert field survey that can be costly, motivating the exploration of AI-driven tools to automate and refine this process. While most AI-driven habitat mapping depends on remote sensing, it is often constrained by sensor availability, weather, and coarse resolution. In contrast, ground-level imagery captures essential structural and compositional cues invisible from above and remains underexplored for robust, fine-grained habitat classification. This study addresses this gap by applying state-of-the-art deep neural network architectures to ground-level habitat imagery. Leveraging data from the UK Countryside Survey covering 18 broad habitat types, we evaluate two families of models - convolutional neural networks (CNNs) and vision transformers (ViTs) - under both supervised and supervised contrastive learning paradigms. Our results demonstrate that ViTs consistently outperform state-of-the-art CNN baselines on key classification metrics (Top-3 accuracy = 91%, MCC = 0.66) and offer more interpretable scene understanding tailored to ground-level images. Moreover, supervised contrastive learning significantly reduces misclassification rates among visually similar habitats (e.g., Improved vs. Neutral Grassland), driven by a more discriminative embedding space. Finally, our best model performs on par with experienced ecological experts in habitat classification from images, underscoring the promise of expert-level automated assessment. By integrating advanced AI with ecological expertise, this research establishes a scalable, cost-effective framework for ground-level habitat monitoring to accelerate biodiversity conservation and inform land-use decisions at the national scale.
@@ 1 Introduction
Habitats are generally classified according to underlying biophysical characteristics of the land on which they occur and the associated species assemblages. Accurate habitat classification is essential for biodiversity conservation and effective land management. Habitat extent and condition data underpin policy frameworks—such as UK’s Biodiversity Net Gain (zu Ermgassen et al., 2021; GOV.UK, 2023) and EU Biodiversity Strategy for 2030 (European Commission, )—and guide conservation priorities, yet they currently depend on expert field surveys which can be costly or difficult to access, for example for landowners looking to understand their land and its value in relation offsetting or carbon credits (Schofield et al., 2024). Recent advances in artificial intelligence (AI) have demonstrated strong potential to automate image-based ecological classification tasks, making habitat identification more accessible for non-experts, thereby reducing costs and alleviating professional workloads.
@@
To date, most AI-driven habitat mapping works rely on remote sensing data (e.g., satellite or drone imagery) (Van An et al., 2023; Norris et al., 2024), which can be hindered by sensor availability, weather conditions, and relatively coarse spatial resolution (satellite imagery typically several m/pixel) (Díaz-Ireland et al., 2024; Morueta-Holme et al., 2024; Marjani et al., 2025). Ground-level imagery, by contrast, captures local structural and compositional cue for habitats (Perrett et al., 2023), which are invisible from above but critical for precise habitat classification. While ground-level images are more readily captured by consumer-grade devices as opposed to a complicated remote sensing pipeline, this avenue remains largely underexplored for developing more comprehensive and robust AI models for habitat recognition (Mall et al., ; Morueta-Holme et al., 2024). In this context, we approach automatic habitat classification using ground-level imagery. By leveraging state-of-the-art AI techniques on richly annotated imagery, we seek to complement existing survey methods, to support cost-effective field-level habitat monitoring and land cover mapping.
@@
This work aims to develop a model that is capable of automatically classifying diverse habitats based on ground-level imagery and with the long-term vision of deploying it on end-user devices in real-world settings (deployment itself falls beyond the scope of this paper). We leverage deep learning approaches, which have become dominant in various computer vision applications, to achieve this goal. A deep neural network is trained on thousands of labelled ground-level images provided by UK Centre for Ecology & Hydrology (UKCEH) Countryside Survey (CS) (Carey et al., 2008), which represents a wide variety of environments across the UK. Based on the UK Habitat (UKHab) Classification (UKHab Ltd, 2023) with a hierarchical habitat system, each image is annotated by experts with one of the level 3 (L3) habitat subcategories–fine-grained subdivisions of level 2 (L2) habitats (for example, grassland at L2 is split into types such as acid grassland, improved grassland, etc.). Appendix C provides detailed descriptions of L2 and L3 habitats as defined in the UKHab.
@@
  • Model architectures. Convolutional Neural Networks (CNNs) are adopted as our baseline, reflecting the prevailing deep learning model in image classification. We further study Vision Transformer (ViT) (Dosovitskiy et al., 2020), whose self-attention mechanism (Vaswani et al., 2017) can more naturally capture global context in ground-level photographs.
  • Training paradigms. Supervised learning with a standard cross-entropy objective is first employed for the multiclass classification problem, to establish performance baselines for each architecture. We then investigate Supervised Contrastive Learning (SupCon) (Khosla et al., 2020) to pretrain the image encoder, aiming to learn tighter, more separable representations for visually similar habitats.
With respect to the classification performance, ViT outperforms stat-of-the-art CNNs consistently by noticeable margins in our experiments. GradCAM (Selvaraju et al., 2017), a model interpretability technique, reveals that ViT makes predictions based on broader and more consistent visual features and avoiding distracting objectives. For the training paradigms, SupCon is found to further improve the performance of ViT by separating visually similar habitats in the embedding space. Finally, we test our best model against human experts, demonstrating that ViT can match and even outperform three highly knowledgeable ecologists for image interpretation.
@@
  • We introduce a novel ground-level habitat classification framework that leverages real-world imagery to address a gap in tools requiring habitat classification (including carbon footprinting, biodiversity offsetting, etc.) for non-experts. Our approach expands the ecological impact of AI by providing more accessible habitat assessments, allowing for extensive habitat classification across the country.
  • Our comparison between state-of-the-art CNNs and ViTs for the habitat classification task demonstrates that ViT is a more suitable deep neural network architecture than CNN for ground-level habitat classification with improved classification performance and interpretability.
  • Our investigation of supervised learning and SupCon demonstrates the latter training paradigm is beneficial for the model to distinguish visually similar habitat. The application of SupCon can boost the classification reliability in real world scenarios.
  • We compare our best model with three experienced human experts in interpreting habitats from a representative subset of level 3 habitat photographs. Our model performs on par with the top expert. This study validates our approach, reveals the remaining challenges, and directs future work toward fully realising the potential of AI-based ground-level habitat classification.
2 Related Work
Computer Vision for Ecology.
Data-driven computer vision has become instrumental in ecological and environmental research by leveraging the growing volume of imagery from diverse sensors to deliver social and economic benefits (Reynolds et al., 2025; Oba and Doi, 2025). Applications include species detection and monitoring, where deep learning models automate identification and population assessment from camera traps (Ratnayake et al., 2021; Bjerge et al., 2022; Xu et al., 2024); underwater and benthic habitat surveys (Zhong et al., 2023), which leverage images recorded by underwater cameras to monitor coral reef growth; plant recognition (Joly et al., 2024; Wang, 2025), employing large public datasets to train deep neural networks to classify plant species (e.g. PlantCEF (Pl@ntNet Consortium, 2025)); and habitat classification (Van An et al., 2023; Norris et al., 2024), largely relying on remote sensing imagery to classify natural habitat types. Among these domains, habitat classification is most closely aligned with our work. Accordingly, we focus our background review on this application—examining the data sources and model architectures where research gaps emerge.
Data Sources for Habitat Classification.
Most data-driven habitat classification studies rely on remote sensing imagery, including satellite data (Praticò et al., 2021; Sittaro et al., 2022; Marcinkowska-Ochtyra et al., 2023; Van An et al., 2023) and drone-based images (Retallack et al., 2022; Norris et al., 2024). While these sources enable coverage of large areas, they depend on favorable environmental conditions, require specialised devices and expertise, and focus on a single top-down perspective, which limits their practicality for non-experts seeking timely, user-friendly, and precise habitat evaluations (Morueta-Holme et al., 2024). In water environments, images captured by underwater camera traps may be used for underwater and benthic habitat surveys (Yasir et al., 2021). On land, ground-level imagery can be easily captured by on-site surveyors or even citizen scientists using consumer devices, offering a more cost-effective and flexible alternative for detailed habitat classification. For example, DeepVerge (Perrett et al., 2023) demonstrates that street-view images can be used to train CNNs for counting grassland species. Saadeldin et al. (2022) shows that a simple CNN trained on a limited number of ground-level imagery achieves accurate classifications over three categories of grassland management intensity. Martinez-Sanchez et al. (2024) applies semantic segmentation to decompose visual features in ground-level photos and uses these features to classify coarse habitat classes (e.g., grassland, cropland), equivalent to UKHab Level 2 categories. However, despite this potential, ground-level imagery remains underexplored to address habitat classification across a variety of subcategories (e.g. different grasslands, woodlands, wetlands).
Models for Habitat Classification
Previous studies in image-based habitat classification (Gómez-Ríos et al., 2019; Praticò et al., 2021; Sittaro et al., 2022; Marcinkowska-Ochtyra et al., 2023; Van An et al., 2023) have predominantly used classical machine learning models, such as support vector machines (Cortes and Vapnik, 1995) and random forests (Breiman, 2001), as well as deep neural networks, specifically CNNs, such as VGG (Simonyan and Zisserman, 2014) and ResNet (He et al., 2016). While these models are effective, they can struggle to address the increasing challenges of learning from more complex images and the growing demand for generalising to a wide range of habitat types. In the family of deep neural networks, the vision transformer (ViT) (Dosovitskiy et al., 2020) has emerged as a state-of-the-art alternative to CNNs in various computer vision tasks. Compared to the CNN that only extracts isolated regional features, the self-attention mechanism (Vaswani et al., 2017) - the foundation of ViT - allows this model to capture relationships between visual features across the entire image, leading to superior performance and interpretability in many areas of application. Recent habitat-focused studies incorporating ViTs (Jamali et al., 2023; Díaz-Ireland et al., 2024; Marjani et al., 2025) have shown promising results, but remain limited to a narrow set of habitats (e.g., grassland, wetland) and typically use satellite imagery rather than ground-based photographs. Consequently, there is a clear opportunity to adapt the ViT to more comprehensive, on-the-ground habitat classification.
Summary.
This work addresses two clear research gaps by: 1) Emphasising habitat classification based on ground-level imagery, which is more flexible for non-experts, and can provides a complementary view to remote sensing 2) Harnessing cutting-edge deep learning architectures to tackle the complexity of diverse habitat types. By combining an underexplored yet highly valuable data source (ground-level images labelled with habitats) with state-of-the-art deep neural networks and training methods, we demonstrate the practical applicability of AI-driven habitat monitoring.
@@ 3.1 Data
Our study focuses on the habitats of the United Kingdom. The habitat images come from the UK Countryside Survey (CS) which covers a randomly stratified sample of UK countryside (Wood et al., 2017), meaning that habitats are sampled in proportion to their occurrence in the wider countryside. Images used were from the most recent rolling survey (CS2019-2023), which covered 500 1km squares across 5 years, during which between 10 and 16 individuals (in each year) were employed as field surveyors. Around half of field surveyors are permanent staff members at UKCEH with other expert surveyors brought in as required. Field surveyors record species within vegetation plots between May and September of the survey year and take photographs of the plots within the context of the habitat in which they are placed, primarily for plot relocation purposes. Plots are assigned to habitats by the field surveyors according to the Broad and Priority Habitat classifications of Jackson (2000) (UK Biodiversity Group, 2008). Broad Habitats comprise part of the hierarchical category system described in the UK Habitat Classification (UKHab) (UKHab Ltd, 2023).
@@
Initially, the data was cleaned by removing unlabelled habitat images, resulting in a dataset with a total number of 5598 samples, including 17 Broad L3 classes and one L4 class (Bracken), corresponding to 9 coarse L2 classes, as shown in Figure 1. While Bracken is technically defined as an L4 habitat in UKHab, it is also classified as a Broad Habitat and is labelled by field surveyors in the CS dataset alongside the L3 habitats. Improved Grassland and Montane are originally Broad Habitats defined in the UK BAP Priority Habitats system (Joint Nature Conservation Committee, n.d.), but they correspond to the scope of L3 habitats in UKHab and are utilised by field surveyors in the CS dataset. For simplicity and consistency throughout this study, we treat Bracken, Improved Grassland, and Montane as equivalent to L3 classes in UKHab, and all references to ’L3 habitats’ in the CS dataset include these three classes unless otherwise specified.
@@ 3.2.1 Model Architecture
The convolutional operation is the core of a CNN for modelling images. In particular, convolutional filters represented by matrices slide across the input image and perform element-wise multiplication and summation to detect visual features, such as edges, shapes, and colours (Goodfellow et al., 2016). The convolutional operation processes the images in local patches. Each time it is calculated, only a small region on the image is activated, whereas the remaining areas are not taken into account. This locality bias assumes that spatially proximate pixels are highly correlated, enabling efficient extraction of meaningful visual features. However, in ground-level habitat imagery–where scenes often contain broad scenes and multiple, potentially distracting objects–this local focus may limit the ability of a CNN to match interpretable visual features to the correct label.
@@
Figure 3: Contrasting locality bias in CNNs with global self-attention in ViT for ground-level habitat classification. A ground-level photograph of a mixed woodland scene is partitioned into a conceptional 3×3 grid of patches. Top Row (CNN Pipeline): A small convolutional filter “slides” over each patch in turn, but never directly integrates information from non-adjacent patches. The resulting GradCam shows concentrated activations only in the immediate neighbourhood of high-contrast features (tree trunks), but misses contextual cues in the background undergrowth. Bottom Row (ViT Pipeline): All nine patches are first embedded into vectors. Lines emanate from an exemplary patch to every other patch, denoting pairwise attention, a thinker line indicates a stronger connection. ViT’s attention mechanism fuses information from both near and far patches. The ViT heatmap highlights both prominent tree trunks and the surrounding trunks, indicating that the model has learned to integrate long-range context—essential for distinguishing forest types.

### v2

@@ Abstract
Habitat assessment at local scales — critical for enhancing biodiversity and guiding conservation priorities — often relies on expert field surveys that can be costly, motivating the exploration of AI-driven tools to automate and refine this process. While most AI-driven habitat mapping depends on remote sensing, it is often constrained by sensor availability, weather, and coarse resolution. In contrast, ground-level imagery captures essential structural and compositional cues invisible from above and remains underexplored for robust, fine-grained habitat classification. This study addresses this gap by applying state-of-the-art deep neural network architectures to ground-level habitat imagery. Leveraging data from the UK Countryside Survey covering 18 broad habitat types, we evaluate two families of models - convolutional neural networks (CNNs) and vision transformers (ViTs) - under both supervised and supervised contrastive learning paradigms. Our results demonstrate that ViTs consistently outperform state-of-the-art CNN baselines on key classification metrics (Top-3 accuracy = 91%, MCC = 0.66) and offer more interpretable scene understanding tailored to ground-level images. Moreover, supervised contrastive learning significantly reduces misclassification rates among visually similar habitats (e.g., Improved vs. Neutral Grassland), driven by a more discriminative embedding space. Finally, our best model performs on par with experienced ecological experts in habitat classification from images, underscoring the promise of expert-level automated assessment. By integrating advanced AI with ecological expertise, this research establishes a scalable, cost-effective framework for ground-level habitat monitoring to accelerate biodiversity conservation and inform land-use decisions at a national scale.
@@ 1 Introduction
Habitats are generally classified according to underlying biophysical characteristics of the land on which they occur and the associated species assemblages. Accurate habitat classification is essential for biodiversity conservation and effective land management. Habitat extent and condition data underpin policy frameworks—such as UK’s Biodiversity Net Gain (zu2021exploring; defra2023bng) and EU Biodiversity Strategy for 2030 (ec2020biodiv2030)—and guide conservation priorities, yet they currently depend on expert field surveys which can be costly or difficult to access, for example for landowners looking to understand their land and its value in relation offsetting or carbon credits (schofield2024aren). Recent advances in artificial intelligence (AI) have demonstrated strong potential to automate image-based ecological classification tasks, making habitat identification more accessible for non-experts, thereby reducing costs and alleviating professional workloads.
@@
To date, most AI-driven habitat mapping works rely on remote sensing data (e.g., satellite or drone imagery) (van2023high; norris2024comparing). Though remote-sensing-based habitat classification excels at quantifying spatial extent, it can be hindered by sensor availability, weather conditions, and relatively coarse spatial resolution (satellite imagery typically several m/pixel) (diaz2024classification; morueta2024unlocking; marjani2025novel). Ground-level imagery, by contrast, captures local structural and compositional cues for habitats (perrett2023deepverge), which are invisible from above but critical for precise habitat classification. It is effective for site-scale habitat assessment and species identification and can be collected with consumer devices rather than a complex remote-sensing pipeline. However, this approach has been less studied for AI-driven habitat classification. In the UK, image analysis is widely used to help both expert and non-expert field surveyors identify species, but no such equivalent currently exists for habitats. Yet, habitat classification is important for mapping and decision-making regarding land use and management, as well as in the context of BNG, a government supported approach currently in use in England to ensure that new projects deliver measurable improvement in nature either on site, or in areas away from the development site (offsets). (mall2312remote; morueta2024unlocking) In this context, we approach automatic habitat classification using ground-level imagery. By leveraging state-of-the-art AI techniques on richly annotated images from the UK Countryside Survey (carey2008countryside), we seek to support cost-effective site-level habitat monitoring, land cover mapping, and complement established approaches such as remote sensing.
Recently, data-driven computer vision has become instrumental in ecological and environmental research by leveraging the growing volume of imagery from diverse sensors to deliver social and economic benefits (reynolds2025potential; oba2025accelerating). Applications include species detection and monitoring, where deep learning models automate identification and population assessment from camera traps (ratnayake2021tracking; bjerge2022real; xu2024advancing); underwater and benthic habitat surveys (zhong2023fine), which leverage images recorded by underwater cameras to monitor coral reef growth; plant recognition (joly2024overview; wang2025plant), employing large public datasets to train deep neural networks to classify plant species (e.g. PlantCEF (plantnet2025)); and habitat classification (van2023high; norris2024comparing).
Within these applications, habitat classification is most directly aligned with our work. We therefore review it in more detail, focusing on data sources and model architectures where research gaps emerge. Currently, habitat classification remains particularly reliant on remote sensing, including satellite data (pratico2021machine; sittaro2022machine; marcinkowska2023natura; van2023high) and drone-based images (retallack2022using; norris2024comparing). While these sources enable coverage of large areas, they depend on favourable environmental conditions, require specialised devices and expertise, and focus on a single top-down perspective, which limits their practicality for non-experts seeking timely, user-friendly, and precise habitat evaluations (morueta2024unlocking). Compared to satellite imagery, ground-level imagery can be easily captured by on-site surveyors or even citizen scientists using consumer devices, offering a cost-effective alternative for detailed habitat classification. For example, DeepVerge (perrett2023deepverge) demonstrates that street-view images can be used to train CNNs for counting grassland species. saadeldin2022using shows that a simple CNN trained on a limited number of ground-level imagery achieves accurate classifications over three categories of grassland management intensity. martinez2024automatic applies semantic segmentation to decompose visual features in ground-level photos and uses these features to classify coarse habitat classes (e.g., grassland, cropland), equivalent to UKHab Level 2 categories. However, despite this potential, ground-level imagery remains underexplored to address habitat classification across a variety of subcategories (e.g. different grasslands, woodlands, wetlands).
Regarding the adopted models for habitat classification, previous studies (gomez2019towards; pratico2021machine; sittaro2022machine; marcinkowska2023natura; van2023high) have predominantly used classical machine learning models, such as support vector machines (cortes1995support) and random forests (breiman2001random), as well as deep neural networks, specifically CNNs, such as VGG (simonyan2014very) and ResNet (he2016deep). While these models are effective, they can struggle to address the increasing challenges of learning from more complex images and the growing demand for generalising to a wide range of habitat types. Conversely, the Vision transformer (ViT) (dosovitskiy2020image) has emerged as a state-of-the-art alternative to CNNs in various computer vision tasks. Compared to the CNN that only extracts isolated regional features, the self-attention mechanism (vaswani2017attention) - the foundation of ViT - allows this model to capture relationships between visual features across the entire image, leading to superior performance and interpretability in many applications. Recent habitat-focused studies incorporating ViTs (jamali2023wetmapformer; diaz2024classification; marjani2025novel) have shown promising results, but remain limited to a narrow set of habitats (e.g., grassland, wetland) and typically use satellite imagery rather than ground-level photographs. Consequently, there is a clear opportunity to adapt the ViT to more comprehensive, ground-level habitat classification.
@@
This work aims to develop a model that is capable of automatically classifying diverse habitats based on ground-level imagery and with the long-term vision of deploying it on end-user devices in real-world settings (deployment itself falls beyond the scope of this paper). We leverage deep learning approaches, which have become dominant in various computer vision applications, to achieve this goal. A deep neural network is trained on thousands of labelled ground-level images provided by UK Centre for Ecology & Hydrology (UKCEH) Countryside Survey (CS) (carey2008countryside), which represents a wide variety of environments across the UK. Based on the UK Habitat (UKHab) Classification (UKHab2023) with a hierarchical habitat system, each image is annotated by experts with one of the level 3 (L3) habitat subcategories–fine-grained subdivisions of level 2 (L2) habitats (for example, grassland at L2 is split into types such as Acid Grassland, Improved Grassland, etc.). Appendix C provides detailed descriptions of L2 and L3 habitats as defined in the UKHab.
@@
  • Model architectures. Convolutional Neural Networks (CNNs) are adopted as our baseline, reflecting the prevailing deep learning model in image classification. We further study Vision Transformer (ViT) (dosovitskiy2020image), whose self-attention mechanism (vaswani2017attention) can more naturally capture global context in ground-level photographs.
  • Training paradigms. Supervised learning with a standard cross-entropy objective is first employed for the multiclass habitat classification problem, to establish performance baselines for each architecture. We then investigate Supervised Contrastive Learning (SupCon) (khosla2020supervised) to pretrain the image encoder, aiming to learn tighter, more separable representations for visually similar habitats.
@@
  • Ground-level habitat classification benchmark. Our benchmark leverages ground-level imagery to address a gap in tools requiring habitat classification (including carbon footprinting, biodiversity offsetting, etc.) for non-experts. It expands the ecological impact of AI by providing more accessible habitat assessments, allowing for extensive habitat classification across the country.
  • Model architecture finding. Our systematic comparison of representative CNNs and ViTs demonstrates that ViTs consistently outperform comparable CNNs under the a broad range of evaluation metrics for habitat classification.
  • Training paradigm investigation. We investigate a two-stage supervised contrastive (SupCon) learning paradigm (contrastive pretraining and classifier fine-tuning), which delivers modest gains over classical supervised learning for our best-performing model and reduces major confusions (e.g., Improved and Neutral Grassland), thereby improving reliability for real-world deployment.
  • Human expert comparison. We compare our best model with three experienced ecologists. On a stratified subset, ViT combining with SupCon is competitive with the best human expert in interpreting habitats from level 3 habitat photographs, while revealing limits and opportunities for our approach.
@@ 2.1 Data
Our study focuses on the habitats of the United Kingdom. The habitat images come from the UK Countryside Survey (CS) which covers a randomly stratified sample of UK countryside (wood2017long), meaning that habitats are sampled in proportion to their occurrence in the wider countryside. Images used were from the most recent rolling survey (CS2019-2023), which covered 500 1km squares across 5 years, during which between 10 and 16 individuals (in each year) were employed as field surveyors. Around half of field surveyors are permanent staff members at UKCEH with other expert surveyors brought in as required. Field surveyors record species within vegetation plots between May and September of the survey year and take photographs of the plots within the context of the habitat in which they are placed, primarily for plot relocation purposes. Specifically, the surveyor is advised to record the location of the plot by taking several ground-level photographs at a near-earth height (1–2 m above ground) using smartphones, including plot markers in the foreground, the wider context of the plot and any significant (easily refindable and preferably permanent features) in the background where possible. These photographs depict the habitat context in the plot. Plots are assigned to habitats by the field surveyors according to the Broad and Priority Habitat classifications of Jackson (2000) (jncc_bap_descriptions). Broad Habitats comprise part of the hierarchical category system described in the UK Habitat Classification (UKHab) (UKHab2023).
@@
Initially, the data was cleaned by removing unlabelled habitat images, resulting in a dataset with a total number of 5598 samples, including 17 Broad L3 classes and one L4 class (Bracken), corresponding to 9 coarse L2 classes, as shown in Figure 1. While Bracken is technically defined as an L4 habitat in UKHab, it is also classified as a Broad Habitat and is labelled by field surveyors in the CS dataset alongside the L3 habitats. Improved Grassland and Montane are originally Broad Habitats defined in the UK BAP Priority Habitats system (jncc_bap_habitats), but they correspond to the scope of L3 habitats in UKHab and are utilised by field surveyors in the CS dataset. For simplicity and consistency throughout this study, we treat Bracken, Improved Grassland, and Montane as equivalent to L3 classes in UKHab, and all references to ’L3 habitats’ in the CS dataset include these three classes unless otherwise specified.
@@ 2.2.1 Model Architecture
The convolutional operation is the core of a CNN for modelling images. In particular, convolutional filters represented by matrices slide across the input image and perform element-wise multiplication and summation to detect visual features, such as edges, shapes, and colours (goodfellow2016deep). The convolutional operation processes the images in local patches. Each time it is calculated, only a small region on the image is activated, whereas the remaining areas are not taken into account. This locality bias assumes that spatially proximate pixels are highly correlated, enabling efficient extraction of meaningful visual features. However, in ground-level habitat imagery–where scenes often contain broad scenes and multiple, potentially distracting objects–this local focus may limit the ability of a CNN to match interpretable visual features to the correct label.
@@
Figure 3: The conceptual difference between how CNN and ViT process the same ground-level habitat image. A ground-level photograph of a mixed woodland scene is partitioned into a conceptional 3×3 grid of patches. In the top row, the CNN extracts visual features using a sliding filter that only sees one patch at a time, making it tend to focus on prominent features in local regions. This locality is reflected in the GradCAM heatmap, where the model concentrates on the trees in the central region (heated area). In contrast, the ViT (bottom row) allows each patch to “attend” to every other patch through self-attention links. The lines in the diagram are a schematic way to show this idea—the thicker lines simply indicate stronger relationships between patches. The ViT therefore considers the global context of the scene, linking near and far regions. The GradCAM heatmap of ViT highlights areas not only on the central trees but also those along the edges, reflecting a broader contextual understanding. This behaviour aligns more closely with human interpretation of ground-level habitat photos, where habitat context often spans the entire scene rather than a single region.